1. Audio cassette tapes of a particular brand are claimed by the manufacturer to give, on average, at least 60 minutes of playing time.  After receiving some complaints, the manufacturer’s quality control manager obtains a random sample of 64 tapes and measures the playing time, t minutes, of each.  The results are summarized by

 t=3953.28 and t²=244,557.00


a. Constructing a symmetric 99% confidence interval for the population mean playing time of the brand of tape.

b. Does the confidence interval support the customer’s complaints? Give a reason for your answer.



1
Expert's answer
2021-11-18T17:35:13-0500

a)

average of sample:

"\overline{x}=3953.28/64=61.77"


"z=\frac{\overline{x}-\mu}{\sigma/\sqrt n}"


"\sigma=\sqrt{t^2/(n-1)}=\sqrt{244557/63}=62"

critical values for 99% confidence interval:

"z=\pm 2.576"


"-2.576<\frac{61.77-\mu}{62/\sqrt{64}}<2.576"


"-19.96<61.77-\mu<19.96"


"41.81<\mu<81.73"


b.

The confidence interval support the customer’s complaints, because mean playing time can be less than 60 minutes (minimal value is 41.81 min).